I have an algebra problem. What it is is a letter from a basic math student and I need to come up with the equation and the answer then be able to explain it to them in a simple way they can understand.
Here is the basic math students question.. During his shift at the school carnival he sold 112 drinks. He thought he only needed to keep track of the total amount of drinks sold, but it turns out they want to know how many bottles of water and how many bottles of soda were sold as separate numbers. The question is: If the number of bottles of water sold was 4 less than 3 times the number of bottles of soda, how many of each did he sell ?
I have to figure out the answer ( how many water and how many sodas) explain how i set up the equation and solved it. So in case the situation happens again they could answer the question.

Answer by ankor@dixie-net.com(22740)

During his shift at the school carnival he sold 112 drinks.
He thought he only needed to keep track of the total amount of drinks
sold, but it turns out they want to know how many bottles of water
and how many bottles of soda were sold as separate numbers.
The question is: If the number of bottles of water sold was 4 less
than 3 times the number of bottles of soda, how many of each did he sell?
:
Let w = no. of bottles of water sold
Let s = no. of bottles of soda
:
Total sold equation
w + s = 112
w = (112-s); use this for substitution
:
"the number of bottles of water sold was 4 less
than 3 times the number of bottles of soda,"
w = 3s - 4
A big mistake here. Let me correct this
Replace w with (112-s)
112-s = 3s - 4
112 + 4 = 3s + s
116 = 4s
s = 116/4
s = 29 sodas
then
w = 112 - 29
w = 83 waters sold
:
See if that checks out in the statement:
"the number of bottles of water sold was 4 less
than 3 times the number of bottles of soda,
83 = 3(29) - 4
83 = 87 - 4